# On Jiang’s wavefront sets conjecture for representations in local Arthur packets

### 13 Mar 2026 v1

@@ 5. Construction of elements in local Arthur packets
$\dim_{\hat{\mathfrak{g}}_{1}^{\prime}}({\mathcal{O}}_{1})+\dim_{\hat{\mathfrak{g}}_{2}^{\prime}}({\mathcal{O}}_{2})>\dim_{\hat{\mathfrak{g}}_{1}^{\prime}}((\underline{p}(\psi_{1})^{t})_{\widehat{\mathrm{G}}_{1}^{\prime}})+\dim_{\hat{\mathfrak{g}}_{2}^{\prime}}((\underline{p}(\psi_{2})^{t})_{\widehat{\mathrm{G}}_{2}^{\prime}}).$
Hence, we have either
$\dim_{\hat{\mathfrak{g}}_{1}^{\prime}}({\mathcal{O}}_{1})>\dim_{\hat{\mathfrak{g}}_{1}^{\prime}}((\underline{p}(\psi_{1})^{t})_{\widehat{\mathrm{G}}_{1}^{\prime}}),$
or,
$\dim_{\hat{\mathfrak{g}}_{2}^{\prime}}({\mathcal{O}}_{2})>\dim_{\hat{\mathfrak{g}}_{2}^{\prime}}((\underline{p}(\psi_{2})^{t})_{\widehat{\mathrm{G}}_{2}^{\prime}}).$
That is, either ${\mathcal{O}}_{1}$ is not related to or bigger than $((\underline{p}(\psi_{1})^{t})_{\widehat{\mathrm{G}}_{1}^{\prime}}$, or, ${\mathcal{O}}_{2}$ is not related to or bigger than $(\underline{p}(\psi_{2})^{t})_{\widehat{\mathrm{G}}_{2}^{\prime}}$, which contradicts Conjecture 1.2.
Therefore, Conjecture 1.7 Part (2) is valid. This completes the proof of Theorem 6.4. ∎
Theorem 1.13 follows from similar arguments as in Theorem 6.4. Details are omitted.
Next, we prove Part (2) of Theorem 1.9.
@@
Assume that $\psi=\bigoplus_{i=1}^{r}\phi_{i}\otimes S_{m_{i}}\otimes S_{n_{i}}$ is non-tempered, i.e., there exists $n_{i}>1$. As before, we assume that $\phi_{i}$ has dimension $k_{i}$, let $a_{i}=k_{i}m_{i}$ and $b_{i}=n_{i}$. Then $\underline{p}_{\psi}=[b_{1}^{a_{1}}\cdots b_{r}^{a_{r}}]$ and by Theorem 1.1, $\mathfrak{p}^{m}({\pi}_{\psi})=\{[b_{1}^{a_{1}}\cdots b_{r}^{a_{r}}]^{t}\}.$ Recall that $I=\{i|b_{i}\text{ odd}\}$, $J=\{i|b_{i}\text{ even}\}$, and $\psi^{1}=\sum_{i\in I}\psi_{i}$, $\psi^{2}=\sum_{j\in J}\psi_{j}$. Since $\psi$ is non-tempered, one can easily see that the two partitions $((\underline{p}(\psi^{1})^{t})^{-})_{{\mathrm{G}}_{1}^{\prime}}$ and $(\underline{p}(\psi_{2})^{t})_{\widehat{\mathrm{G}}_{2}^{\prime}}$ when ${\mathrm{G}}_{n}={\mathrm{Sp}}_{2n}$, or the two partitions $(\underline{p}(\psi_{i})^{t})_{\widehat{\mathrm{G}}_{i}^{\prime}}$, $i=1,2$, when ${\mathrm{G}}_{n}={\mathrm{SO}}_{2n+1},{\mathrm{SO}}_{2n}$, cannot be both parametrizing regular nilpotent orbits.
Assume that there exists a generic representation $\pi_{0}\in\widetilde{\Pi}_{\psi}$. Then there exists a regular nilpotent orbit ${\mathcal{O}}_{0}\in{\mathcal{N}}_{\mathfrak{g}_{n}}$ such that the coefficient $c_{{\mathcal{O}}_{0}}(\pi_{0})$ in the character expansion of $\pi_{0}$ is nonzero, indeed it is $1$ by the result in [MW87] and the uniqueness of Whittaker models. Since for any $\pi\in\widetilde{\Pi}_{\psi}$, $c_{{\mathcal{O}}_{0}}(\pi)$ is either one or zero and $s=1$, by Proposition A.4, we can see that $\sum_{\pi\in\widetilde{\Pi}_{\psi}}\omega_{\pi}(c_{{\mathrm{G}}_{n}})c_{{\mathcal{O}}_{0}}(\pi)\neq 0$.
When ${\mathrm{G}}_{n}={\mathrm{Sp}}_{2n}$, using similar arguments as in the proof of Theorem 6.4, one can see that there exist nilpotent orbits ${\mathcal{O}}_{1}\in{\mathcal{N}}_{{\mathfrak{g}}_{1}^{\prime}}$ and ${\mathcal{O}}_{2}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{2}^{\prime}}$ such that both $c_{{\mathcal{O}}_{1}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{{\mathcal{O}}_{1}}({{}\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}})$ and $c_{{\mathcal{O}}_{2}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{{\mathcal{O}}_{2}}({{}\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}})$ are nonzero, and
$\begin{split}&\,\dim(\mathfrak{g}_{n})-\dim_{\mathfrak{g}_{n}}({\mathcal{O}}_{0})\\ =&\,\dim({\mathfrak{g}}_{1}^{\prime})+\dim(\hat{\mathfrak{g}}_{2}^{\prime})-\dim_{{\mathfrak{g}}_{1}^{\prime}}({\mathcal{O}}_{1})-\dim_{\hat{\mathfrak{g}}_{2}^{\prime}}({\mathcal{O}}_{2}).\end{split}$
@@
$\dim_{{\mathfrak{g}}_{1}^{\prime}}({\mathcal{O}}_{1})+\dim_{\hat{\mathfrak{g}}_{2}^{\prime}}({\mathcal{O}}_{2})=\dim({\mathfrak{g}}_{1}^{\prime})+\dim(\hat{\mathfrak{g}}_{2}^{\prime})-n_{1}-n_{2},$
@@
When ${\mathrm{G}}_{n}={\mathrm{SO}}_{2n+1},{\mathrm{SO}}_{2n}$, similarly, one can see that there exist nilpotent orbits ${\mathcal{O}}_{i}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{i}^{\prime}}$ such that $c_{{\mathcal{O}}_{i}}(\widetilde{\pi}_{\psi^{i}})\hat{\mu}_{{\mathcal{O}}_{i}}({{}\widetilde{f}^{i}_{\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}}})\neq 0$, $i=1,2$, and
$\begin{split}&\,\dim(\mathfrak{g}_{n})-\dim_{\mathfrak{g}_{n}}({\mathcal{O}}_{0})\\ =&\,\dim(\hat{\mathfrak{g}}_{1}^{\prime})+\dim(\hat{\mathfrak{g}}_{2}^{\prime})-\dim_{\hat{\mathfrak{g}}_{1}^{\prime}}({\mathcal{O}}_{1})-\dim_{\hat{\mathfrak{g}}_{2}^{\prime}}({\mathcal{O}}_{2}).\end{split}$
$\dim_{\hat{\mathfrak{g}}_{1}^{\prime}}({\mathcal{O}}_{1})+\dim_{\hat{\mathfrak{g}}_{2}^{\prime}}({\mathcal{O}}_{2})=\dim(\hat{\mathfrak{g}}_{1}^{\prime})+\dim(\hat{\mathfrak{g}}_{2}^{\prime})-n_{1}-n_{2},$
since $n=n_{1}+n_{2}$. Therefore, both ${\mathcal{O}}_{1}$ and ${\mathcal{O}}_{2}$ are regular orbits, which is again a contradiction by Conjecture 1.2 and the discussion above.
@@
At last, we prove Part (3) of Theorem 1.9.
@@
and $n^{*}=\big\lfloor\frac{\sum_{b_{i}\text{ odd }}a_{i}}{2}\big\rfloor$. Assume that Conjecture 1.2 is true. Then Conjecture 1.7 Part (3) holds for the following cases.
@@
    then (6.7) holds and thus Conjecture 1.7 Part (3) is valid.
@@
    then (6.8) holds and thus Conjecture 1.7 Part (3) is valid.
@@
    In particular, if all $b_{i}$ are of the same parity, then (6.9) holds and thus Conjecture 1.7 Part (3) is valid.
@@
To prove Jiang’s Conjecture 1.7 Part (3), we just need to show that for any $\underline{p}>\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi))$, $\underline{p}\notin\mathfrak{p}^{m}(\sigma)$, which is implied by Theorem 6.4 under the validity of Conjecture 1.2 and Proposition A.4. ∎
@@
First, note that with the above assumptions in Cases (1) – (3), $\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi))$ will be exactly the partition parametrizing the subregular nilpotent orbits of $\mathfrak{g}_{n}$. On the other hand, by Lemma 5.5 and Theorem 5.6, we have constructed an irreducible admissible representation $\sigma\in\widetilde{\Pi}_{\psi}$ such that $\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi))\in\mathfrak{p}(\sigma)$. Hence, it suffices to show that there is no generic element in the local Arthur packet corresponding to $\psi$. Since $\psi$ is a non-tempered local Arthur parameter, this is clear from Conjecture 1.6. ∎

### 06 May 2026 v2

@@ 5. Construction of elements in local Arthur packets
$\begin{split}\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{\mathcal{O}}((f)_{t^{2}})=|t|^{\dim(\mathfrak{g}_{n})-\dim(\mathfrak{g}_{n}^{\theta})}c_{{\mathcal{O}}_{0}}(\widetilde{\pi}_{\psi})\hat{\mu}_{{\mathcal{O}}_{0}}(({{}\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{n}}}})_{t^{2}}).\end{split}$
On the other hand,
$\hat{\mu}_{\mathcal{O}}((f)_{t^{2}})$ $=|t|^{2\dim(\mathfrak{g_{n}})-\dim({\mathcal{O}})}\hat{\mu}_{\mathcal{O}}(f),$
$\hat{\mu}_{{\mathcal{O}}_{0}}(({{}\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{n}}}})_{t^{2}})$ $=|t|^{2\dim(\mathfrak{g}_{n}^{\theta})-\dim({\mathcal{O}}_{0})}\hat{\mu}_{{\mathcal{O}}_{0}}({{}\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{n}}}}).$
Therefore, there exists ${\mathcal{O}}_{1}\in{\mathcal{N}}_{\mathfrak{g}_{n}}$ such that $c_{{\mathcal{O}}_{1}}(\widetilde{\Pi}_{\psi})\neq 0$ and
$2\dim(\mathfrak{g_{n}})-\dim({\mathcal{O}}_{1})=\dim(\mathfrak{g}_{n})-\dim(\mathfrak{g}_{n}^{\theta})+2\dim(\mathfrak{g}_{n}^{\theta})-\dim({\mathcal{O}}_{0}).$
Hence, when ${\mathrm{G}}_{n}={\mathrm{SO}}_{2n+1},{\mathrm{Sp}}_{2n}$, $\dim({\mathcal{O}}_{1})=\dim({\mathcal{O}}_{0})>\dim_{\mathfrak{g}_{n}^{\theta}}((\underline{p}(\psi)^{t})_{{\mathrm{G}}_{n}^{\theta}})$ which is equal to
$\dim_{\mathfrak{g}_{n}}(\eta_{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}([b_{1}^{a_{1}}\cdots b_{r}^{a_{r}}]))$
by Proposition 4.2, contradicting our assumption. When ${\mathrm{G}}_{n}={\mathrm{SO}}_{2n}$, $\dim({\mathcal{O}}_{1})=\dim(\mathfrak{g_{n}})-\dim(\mathfrak{g}_{n}^{\theta})+\dim({\mathcal{O}}_{0})>\dim(\mathfrak{g_{n}})-\dim(\mathfrak{g}_{n}^{\theta})+\dim_{\mathfrak{g}_{n}^{\theta}}((\underline{p}(\psi)^{t})_{{\mathrm{G}}_{n}^{\theta}})=\dim_{\mathfrak{g}_{n}}(\eta_{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}([b_{1}^{a_{1}}\cdots b_{r}^{a_{r}}]))$, by Lemma 4.5, also contradicting our assumption.
This completes the proof of Theorem 6.4. ∎
Next, we prove Theorem 1.9(3-a).
@@
Assume that $\psi=\bigoplus_{i=1}^{r}\phi_{i}\otimes S_{m_{i}}\otimes S_{n_{i}}$ is non-tempered, i.e., there exists $n_{i}>1$. As before, we assume that $\phi_{i}$ has dimension $k_{i}$, let $a_{i}=k_{i}m_{i}$ and $b_{i}=n_{i}$. Then $\underline{p}_{\psi}=[b_{1}^{a_{1}}\cdots b_{r}^{a_{r}}]$ and by Theorem 1.1, $\mathfrak{p}^{m}({\pi}_{\psi})=\{[b_{1}^{a_{1}}\cdots b_{r}^{a_{r}}]^{t}\}.$ Recall that $I=\{i|b_{i}\text{ odd}\}$, $J=\{i|b_{i}\text{ even}\}$, and $\psi^{1}=\sum_{i\in I}\psi_{i}$, $\psi^{2}=\sum_{j\in J}\psi_{j}$. Since $\psi$ is non-tempered, one can easily see that the two partitions $(\underline{p}(\psi^{1})^{t})_{{{\mathrm{G}}_{1}^{\prime}}^{\theta}}$ and $(\underline{p}(\psi_{2})^{t})_{{{\mathrm{G}}_{2}^{\prime}}^{\theta}}$ cannot be both parametrizing regular nilpotent orbits.
Assume that there exists a generic representation $\pi_{0}\in\widetilde{\Pi}_{\psi}$. Then there exists a regular nilpotent orbit ${\mathcal{O}}_{0}\in{\mathcal{N}}_{\mathfrak{g}_{n}}$ such that the coefficient $c_{{\mathcal{O}}_{0}}(\pi_{0})$ in the character expansion of $\pi_{0}$ is nonzero, indeed it is $1$ by the result in [MW87] and the uniqueness of Whittaker models. Since for any $\pi\in\widetilde{\Pi}_{\psi}$, $c_{{\mathcal{O}}_{0}}(\pi)$ is either one or zero and $s=1$, we can see that $\sum_{\pi\in\widetilde{\Pi}_{\psi}}c_{{\mathcal{O}}_{0}}(\pi)\neq 0$.
Applying similar arguments as in the proof of Theorem 6.4, one can see that there exist nilpotent orbits ${\mathcal{O}}_{1}\in{\mathcal{N}}_{\mathfrak{g_{1}^{\prime}}^{\theta}}$ and ${\mathcal{O}}_{2}\in{\mathcal{N}}_{\mathfrak{g_{2}^{\prime}}^{\theta}}$ such that both $c_{{\mathcal{O}}_{1}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{{\mathcal{O}}_{1}}({{}\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}})$ and $c_{{\mathcal{O}}_{2}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{{\mathcal{O}}_{2}}({{}\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}})$ are nonzero, and
[…]
@@
$\dim{{}_{\mathfrak{g_{1}^{\prime}}^{\theta}}}({\mathcal{O}}_{1})+\dim{{}_{\mathfrak{g_{2}^{\prime}}^{\theta}}}({\mathcal{O}}_{2})=\dim({\mathfrak{g_{1}^{\prime}}^{\theta}})+\dim({\mathfrak{g_{2}^{\prime}}^{\theta}})-n_{1}-n_{2},$
@@
At last, we prove Theorem 1.9(3-b).
@@
and $n^{*}=\big\lfloor\frac{\sum_{b_{i}\text{ odd }}a_{i}}{2}\big\rfloor$. Assume that Conjecture 1.2 is true. Then Conjecture 1.7(3) holds for the following cases.
@@
    then (6.7) holds and thus Conjecture 1.7(3) is valid.
@@
    then (6.8) holds and thus Conjecture 1.7(3) is valid.
@@
    In particular, if all $b_{i}$ are of the same parity, then (6.9) holds and thus Conjecture 1.7(3) is valid.
@@
To prove Jiang’s Conjecture 1.7(3), it remains to show that for any $\underline{p}>\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi))$, $\underline{p}\notin\mathfrak{p}^{m}(\sigma)$; this is implied by Theorem 6.4 under the validity of Conjecture 1.2. ∎
@@
First, note that with the above assumptions in Cases (1-3), $\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi))$ will be exactly the partition parametrizing the subregular nilpotent orbits of $\mathfrak{g}_{n}$. On the other hand, by Lemma 5.5 and Theorem 5.6, we have constructed an irreducible admissible representation $\sigma\in\widetilde{\Pi}_{\psi}$ such that $\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi))\in\mathfrak{p}(\sigma)$. Hence, it suffices to show that there is no generic element in the local Arthur packet corresponding to $\psi$. Since $\psi$ is a non-tempered local Arthur parameter, this is clear from Conjecture 1.6. ∎
@@
In this appendix, we show that all representations in local Arthur packets have the same central characters, which has its own interests.